### v3

# Game Reasoning Arena: A Framework and Benchmark for Assessing Reasoning Capabilities of Large Language Models via Game Play

Game Reasoning Arena: A Framework and Benchmark for Assessing Reasoning Capabilities of Large Language Models via Game Play
Lucia Cipolina‑Kun LAION Juelich Supercomputing Center (JSC), Research Center Juelich (FZJ) Marianna Nezhurina LAION Juelich Supercomputing Center (JSC), Research Center Juelich (FZJ) Jenia Jitsev LAION Juelich Supercomputing Center (JSC), Research Center Juelich (FZJ)
(August 18, 2025)
The Game Reasoning Arena library provides a framework for evaluating the decision making abilities of large language models (LLMs) through strategic board games implemented in Google’s OpenSpiel library. The framework enables systematic comparisons between LLM‑based agents and other agents (random, heuristic, reinforcement learning agents, etc.) in various game scenarios by wrapping multiple board and matrix games and supporting different agent types. It integrates API access to models via liteLLM, local model deployment via vLLM, and offers distributed execution through Ray. This paper summarises the library’s structure, key characteristics, and motivation of the repository, highlighting how it contributes to the empirical evaluation of the reasoning of LLM and game‑theoretic behaviour. The complete library documentation can be found at https://game-reasoning-arena.readthedocs.io/ and the library is hosted at https://github.com/SLAMPAI/game_reasoning_arena.
1 Introduction
Recent advances in large language models have spurred interest in evaluating their reasoning and planning abilities beyond standard natural language benchmarks. Strategic games offer a controlled environment where agents must plan, adapt and anticipate opponents’ moves, making them a natural testbed for decision maGame Reasoning Arenaaddresses this need by providing a unified framework for playing and analysing games such as Tic–Tac–Toe, Connect Four, Kuhn Poker and matrix games like Prisoner’s Dilemma. It uses Google‑DeepMind’s OpenSpiel library as the underlying game engine. OpenSpiel is an open‑source collection of environments and algorithms for reinforcement learning and search/planning in games; it supports single‑ and multi‑agent, zero‑sum and general‑sum games with perfect and imperfect information and includes tools to analyse learning dynamics. By building on this engine, the framework exposes a flexible API to configure games, agents and evaluation settings.
The repository is organised into several modules:
  • Game registry and environments. Games are registered via decorators in the loaders module, associating a name with an OpenSpiel game loader and a custom environment class; for example, Tic–Tac–Toe and Connect Four have dedicated environment classes. The registry singleton handles dynamic loading and instantiation of games.
  • Agents and policies. Agents implement a compute_action method and include random, human and LLM agents. The policy manager assigns policies based on configuration and uses an agent registry to map strings to classes. The LLM agent queries a language model using a unified backend and extracts an action and optional reasoning, while the random agent selects uniformly from legal actions.
  • Simulation logic. The simulation loop creates an environment, seeds the game, and then repeatedly queries each agent for an action. It handles both turn‑based and simultaneous games, applies actions to the environment, checks for illegal moves, updates rewards and logs outcomes. Optional parallel execution allows multiple episodes or games to run concurrently on several CPU or GPU resources.
  • Architecture overview. An accompanying architecture document outlines how the registry, environment, agent and backend layers interconnect. At a high level, environments derive from a common OpenSpiel wrapper, agents derive from a base agent class, and backends provide model‑agnostic inference. These modules are orchestrated by simulation utilities that coordinate policy assignment and logging.
Overall, the repository emphasises modularity: researchers can add new games or agents by implementing a new environment or agent class and registering it via decorators.
2 Framework Design
The simulation framework builds on the multi-agent reinforcement-learning (RL) paradigm, which wraps the extensive game catalogue of OpenSpiel in a Gymnasium-like interface. This section describes the game loop, environment design, agent interface and language-model inference back-ends, emphasising how these components support rigorous evaluation of LLM reasoning within strategic games.
2.1 Game loop and environment
The codebase leverages OpenSpiel as the backend for the games emulator. The OpenSpiel API interface acts as the RL environment, where it encapsulates the game state, checks the validity of actions and mediates the transition dynamics. At each step, the OpenSpiel API’s output is wrapped in a text prompt, this is passed to the LLM agent as the state observation. The text prompt includes the current state representation, the set of legal actions and for some games a summary of past moves. This design mirrors Gymnasium’s API and supports both turn-based and simultaneous-move games.
2.2 Agent interface and observation structure
Agents implement a unified compute_action method that receives the prompt as observation and returns an action as a dictionary with two items, the selected action and the reasoning behind the chosen action. The prompt asks the LLM to output its reasoning before selecting an action, to avoid ex-post contamination. The environment returns reward signals calculated by OpenSpiel as dictionaries mapping each player to a scalar value, consistent with multi-agent RL practice.
2.3 Language model inference backends
A key feature of our framework is support for multiple inference backends, enabling experiments with both hosted and locally deployed language models. The LiteLLM backend offers a unified API covering more than a hundred providers, including OpenAI, Anthropic, Google and Groq, and allows cost-effective and fast inference through automatic batching, caching and rate-limit handling. It also enables researchers to mix providers within a single experiment, facilitating controlled comparisons across models. The vLLM backend runs language models locally on a GPU via the vLLM inference engine, offering privacy, deterministic performance and offline evaluation capabilities, albeit requiring local model weights and sufficient GPU memory. By decoupling the agent logic from the underlying back-end, our design allows researchers to plug in any model that conforms to the same interface and to investigate how the choice of model and inference infrastructure influences game-play behaviour.
3 Motivation
Evaluating LLMs in strategic contexts is motivated by the desire to understand whether these models can perform complex planning and reasoning beyond text completion. Board and card games involve sequential decision making, imperfect information and adversarial dynamics, making them a rich domain to test capabilities aligned with reinforcement learning and game theory. Board Game Arena provides several incentives for researchers:
  • Multi‑agent testing. The library supports for LLMs vs random, LLM vs human, LLM vs LLM and self‑play settings, enabling comparative analyses of different models and baselines.
  • Diverse games. By including both perfect information games (Tic-Tac-Toe, Connect Four) and hidden information games (Kuhn Poker), as well as matrix games (Prisoner’s Dilemma, Matching Pennies), the framework tests various aspects of strategic reasoning.
  • Flexible backends and cross‑provider evaluation. Models from Groq, Together AI, Fireworks and other providers can be used alongside local vLLM models; this allows researchers to assess how inference speed, token limit and temperature influence performance.
  • Scalability. Integration with Ray enables parallel simulation across multiple CPUs/GPUs or even SLURM clusters, facilitating large‑scale experiments.
  • Extensibility. The library is designed in a flexible and modular way to add new games or agents. This is essential for investigating novel game scenarios or implementing reinforcement learning agents for comparison.
  • Multi‑agent training. The library is designed following Gymansium and RLLIB paradigms for reinforcement learning. This makes it suitable for users to train their own agent in the provided environment.
From a researcher’s point of view, the framework bridges the gap between language modelling and game‑theoretic evaluation. It allows testing whether LLMs can learn to cooperate in iterated Prisoner’s Dilemma, bluff in Kuhn Poker or plan winning sequences in Connect Four. By logging reasoning strings from LLM responses and recording illegal moves, researchers can study failure modes and prompt designs. Given the rapid development of open and proprietary models, the cross‑backend registry is a pragmatic feature that simplifies experimentation. Overall, the Board Game Arena repository offers a structured and extensible platform to interrogate the strategic capabilities of modern language models.
4 Prompting System Architecture
A central feature of our benchmark is its structured prompting system, which enables consistent, analyzable interactions between language models and diverse board game environments. The design is modular and extensible, reflecting best practices in agent–environment interfaces and prompt engineering for LLMs.
4.1 Observation and Decision Cycle
Each game environment implements a state-to-observation transformation that produces a structured dictionary for each agent. This includes:
  • state_string: A human-readable representation of the current game state.
  • legal_actions: A list of all valid moves the agent can take.
  • prompt: A fully constructed prompt string containing task-specific context.
Agents extract the prompt field, which is passed to an LLM backend. Backend-specific adapters apply any required formatting (e.g., chat templates for dialogue models). The returned string is parsed into an action and accompanying reasoning using structured output constraints.
4.2 Hierarchical Prompt Construction
Prompt generation follows a hierarchical strategy, allowing base prompts to be reused while enabling customization for game-specific contexts.
Base Environment Prompts
The generic OpenSpiel environment provides prompts that include:
  • Game name and player role
  • Move number and board visualization
  • List of legal actions
Specialized Prompts
Environments for games with hidden information or contextual semantics—such as Kuhn Poker—extend the prompt with:
  • Private observations (e.g., the player’s hidden card)
  • Game history (e.g., betting sequence)
  • Interpretable action labels based on state semantics
4.3 Standardized Prompt Augmentation
All prompts are further formatted using a standardized wrapper function that appends:
  • A reasoning directive, instructing the LLM to verbalize its strategy.
  • A JSON schema for the output, specifying fields reasoning and action.
This dual-layer prompt format enables structured logging and consistent response parsing.
Example: Prompt in Kuhn Poker
The following is an example of a prompt generated for a Kuhn Poker game. This prompt includes both private and public information, contextualized action labels, and structured response instructions:
You are Player 0 in the game Kuhn Poker.
Your private card: Jack
This is move number: 2
Betting history: [’Check’]
Total pot size: 2 chips
Your contribution: 1 chips
Available actions:
0: Check (stay in the game without betting)
1: Bet (add a chip to the pot)
What action do you choose? Reply only with ’0’ or ’1’.
First, think through the game strategy and explain your reasoning.
Only after that, decide on the best action to take.
Reply only in the following JSON format:
  ’reasoning’: <str>,
  ’action’: <int>
This format enables the LLM to reason through game dynamics and ensures that the action can be extracted programmatically using standard JSON parsing or regular expressions.
4.4 Backend Integration and Parsing
The system supports both raw-text and chat-based models. When needed, prompts are automatically formatted into structured chat messages using templates tailored to each backend. To handle formatting inconsistencies, model outputs are parsed with regular expressions, allowing structured data to be extracted reliably. Any reasoning text generated by the model is preserved for further analysis and visualization.
4.5 Extensibility for New Environments
To support new games, developers can simply override the _generate_prompt() method in their environment. The framework handles downstream formatting and parsing, simplifying integration. The prompt system thus balances flexibility with enforced structure, allowing scalable and interpretable LLM-agent integration across games.
5 Evaluation and Experimental Design
Our evaluation focuses on measuring the strategic competence and reasoning quality of language models beyond simple win rates. For each game we record per-step rewards, outcomes and the sequence of actions and reasoning strings produced by the agents. From these trajectories we derive metrics such as (i) average and maximum cumulative reward, (ii) decision optimality, measured as the proportion of moves matching the equilibrium or optimal policy, (iii) reasoning length and coherence, captured via simple heuristics on the textual rationales, and (iv) error rates, counting illegal or suboptimal moves.
To ensure statistical significance, we run multiple simulations per configuration and report means and standard errors. When comparing models across games or back-ends, we use paired-sample tests and bootstrap confidence intervals to assess differences in performance.
5.1 Experimental design
The framework supports systematic experimentation across a wide range of configurations, enabling controlled evaluation of agent behaviors under varying game setups. It facilitates large-scale simulations, hyperparameter sweeps, and ablation studies through seamless parallelization. Each experiment is initialized with a consistent random seed to ensure reproducibility, and simulations are executed using the framework introduced in Section 2. To enable scalability, we leverage parallelization via Ray, distributing simulation workloads across available hardware resources. On high-performance computing clusters, experiments are scheduled using SLURM to facilitate efficient large-scale evaluations.
5.2 Analysis tools
Beyond aggregate metrics, we provide tools to inspect agent behaviour. Reasoning strings from the LLM agents are logged alongside actions, allowing qualitative analysis of decision-making and failure modes. Post-game processing scripts summarise common reasoning categories, identify when models refer to game-specific concepts and flag hallucinations or rule violations. These tools facilitate deeper insights into model capabilities and guide prompt engineering and model selection.
5.3 Results
To analyse the reasoning processes of LLMs in board game environments, we propose a categorization framework that labels model-generated justifications according to distinct reasoning patterns as shown in Table 1. These include positional reasoning, which emphasizes spatial control of the board (e.g., references to the ”center column” or ”corner”); opponent modeling, where the model anticipates or reacts to the opponent’s strategy; and blocking, which denotes defensive actions aimed at preventing the opponent’s success. Additional categories include winning logic, which refers to direct strategies for achieving victory (e.g., ”winning move”, ”fork”); heuristic-based reasoning, characterized by general evaluative language (e.g., ”best move”, ”advantageous”); and rule-based reasoning, which appeals to known strategies or game rules. Finally, we include a random/unjustified category for responses lacking clear strategic motivation. By associating these types of reasoning with specific lexical cues, this classification enables a structured analysis of LLM decision making, highlighting patterns, strengths, and potential shortcomings in model reasoning across varied game scenarios.
We apply this classification framework to large-scale gameplay logs from multiple LLMs and games, producing the following multi-perspective analysis.
Table 1: Reasoning types with descriptions and representative words.
Reasoning Type Description Words
Positional Focuses on spatial board control (e.g. “center column”) center column, “center square”, “corner”, “edge”
Opponent modeling Refers to or anticipates opponent strategy “opponent”, “they are trying”, “their strategy”, “their move”
Blocking Refers to countering opponent threats “block”, “prevent”, “stop opponent”, “avoid opponent”, “counter”
Winning logic Direct mentions of creating win conditions “win”, “winning move”, “connect”, “fork”, “threat”, “chance of winning”
Heuristic-based General heuristics like “center is best” “best move”, “most likely”, “advantageous”, “better chance”
Rule-based Based on known strategies (forks, chains) according to, “rule”, “strategy”
Random/unjustified Reasoning lacks strategic justification “random”, “guess”
5.3.1 Temporal Adaptation Across Games (Radar Overview)
Figure 1 summarises temporal reasoning patterns for groq_llama3_70b_8192 across seven games. Each polygon corresponds to a game; axes are reasoning types and radii denote normalised frequencies over the course of play. We observe opponent-focused behaviour in tic_tac_toe and prisoners_dilemma, positional emphasis in connect_four, rule-based regularity in matching_pennies, and a larger uncategorised share in matrix_pd.
Figure 1: Temporal reasoning-type patterns across games for groq_llama3_70b_8192.
5.3.2 Aggregate Per-Game Profiles (Stacked Bars and By-Game Views)
Figures 23 report the overall per-game distribution of reasoning types. Stacked bars offer a compact percentage view for each game; the by-game plots show the same information with separate bars per category.
The 70B model exhibits richer variety and stronger opponent modelling in tic_tac_toe and matching_pennies, positional play in connect_four, and uncategorised traces in the matrix games. The 8B model shows heavier reliance on few categories and more frequent uncategorised segments.
Figure 2: Stacked reasoning percentages by game: groq_llama3_70b_8192.
Figure 3: Stacked reasoning percentages by game: groq_llama3_8b_8192.
Figures 4, 5, and 6 show that stronger models adapt their reasoning to the structural demands of each game. For example, the Kimi-K2-Instruct model relies almost exclusively on opponent modeling in Kuhn Poker and Prisoner’s Dilemma, while exhibiting more diverse strategies in Connect Four and Tic-Tac-Toe. GPT-4, in turn, displays greater adaptability by deploying blocking strategies in Connect Four while relying on opponent modeling in sequential games. GPT-3.5-Turbo, however, exhibits weaker specialization, with many outputs uncategorized and lacking clear strategic grounding. These findings suggest that stronger models not only adapt within games but also calibrate their strategies across games.
Figure 4: Reasoning distribution across games for kimi_k2_instruct. Opponent modeling dominates in simpler games, while more complex board games elicit a broader range of reasoning types.
Figure 5: Reasoning distribution for gpt_4. The model adapts its reasoning to game structure, mixing opponent modeling, blocking, and uncategorized strategies.
Figure 6: Reasoning distribution for gpt_3.5_turbo. Compared to GPT-4, the model shows weaker adaptation and a higher proportion of uncategorized reasoning.
5.3.3 Within-Game Evolution (Same Game, Different Turns)
We bin turns and compute per-bin reasoning proportions to detect shifts in strategy within a single game. Heatmaps emphasise concentration and switching; stacked bars show the mixture per bin. In tic_tac_toe, both models shift towards pure Blocking in the endgame, but their openings differ—one starts with Winning Logic then Opponent Modeling, while the other begins with strong Opponent Modeling and a positional midgame.
Figure 7: Within-game reasoning heatmap (tic_tac_toe): groq_llama_3_1_8b_instant.
Figure 8: Turn-binned proportions (tic_tac_toe): groq_llama_3_1_8b_instant.
Figure 9: Within-game reasoning heatmap (tic_tac_toe): groq_llama3_8b_8192.
Figure 10: Turn-binned proportions (tic_tac_toe): groq_llama3_8b_8192.
5.3.4 Per-Game Pie Charts (Detailed Breakdown Examples)
We include game-specific pies to illustrate the detailed mix behind the aggregates. In tic_tac_toe the instant 8B model balances Blocking (50%), Winning Logic (25%), and Opponent Modeling (25%). The 8B 8192 variant diversifies further, and in kuhn_poker the 70B model emphasises Opponent Modeling with smaller rule-based and uncategorised shares.
Figure 11: Pie breakdown (tic_tac_toe): groq_llama_3.1_8b_instant.
Figure 12: Pie breakdown (tic_tac_toe): groq_llama3_8b_8192.
Figure 13: Pie breakdown (kuhn_poker): groq_llama3_70b_8192.
5.3.5 Entropy Across Agents Within a Game
To quantify cross-model diversity of reasoning over time for a fixed game, we compute the Shannon entropy of the reasoning-type distribution at each turn bin, aggregating across agents. Higher entropy indicates a larger spread across reasoning types, whereas zero entropy reflects unanimity.
Figure 14 reports entropy across agents for tic_tac_toe. The groq_llama3_8b_8192 model shows a clear mid-game peak (greater strategic diversity), followed by a collapse to near-zero in late turns, consistent with convergence to Blocking. The other agents remain near zero throughout.
Figure 14: Entropy by turn across agents (tic_tac_toe). Higher values indicate more diverse reasoning distributions across agents at a given turn; zero denotes unanimity.
5.3.6 Synthesis
Taken together, these views reveal a coherent pattern. First, LLMs adapt their reasoning across games in ways aligned with game structure (radar, stacked bars). Second, they adapt within a single game: openings display planning/anticipation, mid-games emphasise board control, and endgames converge to defensive blocking (heatmaps and turn-stacked bars). Third, model size influences diversity—larger models tend to display richer, more balanced profiles, while smaller ones rely on a narrower set of rationales. Finally, entropy confirms that strategic diversity varies over time and across agents, peaking when the search space is most open and collapsing as terminal constraints dominate.
6 Conclusion and Outlook
We have introduced a board-game codebase and benchmark for evaluating the strategic reasoning of large language models. Building on the multi-agent RL paradigm of OpenSpiel, our framework unifies a diverse set of discrete games within a common interface, supports multiple LLM inference backends and records both actions and reasoning strings. The resulting benchmark enables systematic comparisons across models, games and inference modalities, providing insights into how language models plan, adapt and cooperate in strategic settings.
Our experiments outline a broad suite of evaluation metrics and demonstrate how distributed execution facilitates large-scale benchmarking. Although our focus has been on turn-based board and card games, the modular design readily extends to other game genres and to new agent types. Future work will explore richer reasoning analyses, adaptive prompting strategies and integration with reinforcement-learning training loops. We hope that this benchmark will stimulate further research into the game-theoretic capabilities of language models and their applications in human-AI interaction.
7 Related Research
The growing interest in using games to evaluate large language models has led to several complementary benchmarks.
TextArena  Guertler et al. (2025) is an open‑source collection of more than fifty text‑based games designed to train and evaluate agentic behaviour in LLMs. It supports single‑player, two‑player and multi‑player scenarios and provides an online platform where models and humans can compete, with performance tracked via a real‑time TrueSkill leaderboard.
GameArena Hu et al. (2025b) focuses on reasoning skills. It is a dynamic benchmark in which models play live computer games against human opponents to test deductive and inductive reasoning. The benchmark comprises three games, collects over 2,000 gameplay sessions, and analyses the resulting data to measure fine‑grained reasoning capabilities.
In the realm of board and grid‑based games, Topsakal et al. (2024) introduced an extensible benchmark using games such as Tic–Tac–Toe, Connect Four and Gomoku to compare LLMs. The open‑source simulator runs thousands of matches between models from Anthropic, Google, OpenAI and Meta, generating results in multiple formats and creating a leaderboard for analysis.
GameBench Costarelli et al. (2024) extends evaluation to a broader set of strategic games. It defines a cross‑domain benchmark covering nine strategic board, card and social deduction games that test reasoning along axes such as imperfect information, stochasticity and communication. Evaluations of GPT‑3 and GPT‑4 (with chain‑of‑thought or planning scaffolds) reveal that even the best models often underperform random baselines and fail to approach human-level performance.
Board Game Bench Board Game Bench authors (2025) is an a rena‑based benchmark that pits LLMs head‑to‑head in board games. Despite rapid progress, LLMs remain poor at most board games; success requires rule comprehension, opponent modelling, probability estimation, and long‑term planning. The system records wins/losses and uses a Bayesian Elo‑like rating to rank models.
lmgame‑Bench Hu et al. (2025a) converts real video games into robust evaluation benchmarks. The framework integrates perception, memory and reasoning scaffolds and provides a unified environment across games such as Super Mario Bros., Tetris, and Ace Attorney. Reinforcement‑learning experiments on this suite demonstrate both in‑domain and cross‑domain generalisation gains.
Benchmark Game Set Focus/Skills Tested Citation
TextArena 50+ text games Social reasoning (negotiation, deception, theory of mind); real-time leaderboard Guertler et al. (2025)
GameArena 3 live games, 2,000+ sessions Deductive and inductive reasoning; interactive gameplay with human opponents Hu et al. (2025b)
Grid-based LLM benchmark Tic–Tac–Toe, Connect Four, Gomoku Rule understanding, win-rate analysis, strategic variation by prompt Topsakal et al. (2024)
GameBench 9 board/card/social games Reasoning under imperfect info, stochasticity, communication; evaluation with GPT-3/4 Costarelli et al. (2024)
Board Game Bench Head-to-head board games Opponent modelling, long-term planning, Bayesian Elo ranking Board Game Bench authors (2025)
lmgame‑Bench Real video games Vision, memory, planning, and generalisation under contamination-robust setup Hu et al. (2025a)
Table 2: Comparison of recent game-based LLM evaluation benchmarks.
Collectively, these benchmarks indicate that while LLMs excel at many language tasks, strategic gameplay remains challenging. They highlight the importance of dynamic, multi‑agent and interactive environments for evaluating reasoning, planning and social skills, and suggest motivation for frameworks such as Game Reasoning Arena
Appendix A Appendix
The architecture of our codebase is organized into modular components that separate concerns across environment design, agent control, and simulation execution. As illustrated in Figure 15, the simulation process begins with runner.py, which serves as the main entry point and initializes the overall configuration. This script delegates the core simulation logic to simulate.py, which orchestrates the interaction between the environment, game registry, and agent policies. On the left-hand side of the diagram, simulate.py interfaces with registry.py, a module responsible for loading and registering matrix games through the Matrix Game Loader. At the center of the architecture, environment classes such as MatrixGameEnv are defined to implement the game logic and extend a common OpenSpielEnv base class to ensure compatibility with the OpenSpiel interface. On the right-hand side, agent configuration and management are handled by policy_manager.py, which supports multiple agent types, including large language models (LLMs), random agents, and human-controlled agents. Each agent computes its actions based on the current game state and proceeds through the simulation loop via the action computation and game step execution phases. Finally, game outcomes and agent behaviors are recorded through a structured logging mechanism, enabling detailed analysis and post-hoc evaluation. This modular and extensible design facilitates efficient benchmarking of LLM strategies in matrix game settings.
Figure 15: High-level architecture of the Game Reasoning Arena framework for matrix games.
The interaction flow of a single matrix game episode within the Game Reasoning Arena is depicted in Figure 16. The simulation begins with environment initialization via env.reset(), followed by the construction of player-specific observations using the _state_to_observation() method. The simulation loop then enters the Action Computation Phase, where each agent selects an action based on its observation. For Player 0 (an LLM-controlled agent), this involves querying a language model using a prompt that includes the legal actions. The LLM returns both an action and its reasoning (e.g., “Paper beats Rock”). For Player 1 (a random agent), actions are sampled uniformly. In the Simultaneous Step Phase, the environment executes the joint action using env.step(), which internally calls state.apply_actions() to apply the moves and computes rewards via _compute_reward(). After each step, the environment returns updated observations, rewards, and a termination flag. Finally, during the Logging Phase, the system records all relevant information, including the LLM’s reasoning, action choices, and outcome metrics, using structured logs (e.g., SQLiteLogger) and real-time visualization tools such as TensorBoard. This loop continues until the game reaches a terminal state.
Figure 16: Sequence diagram of a single matrix game episode in the Game Reasoning Arena. It illustrates environment setup, LLM and random agent decision-making, game state updates, and structured logging.